OpenStudy (anonymous):

solve for v: u=3y-6v

OpenStudy (compassionate):

Do you know the first step?

OpenStudy (anonymous):

not really

OpenStudy (compassionate):

First you want to divide 3 and add y

OpenStudy (anonymous):

3yu=6?

OpenStudy (compassionate):

Go to MathAway and then type it in there. You'll get the answers.

OpenStudy (shane_b):

\[u=3y-6v\] Subtract 3y from both sides: \[u-3y=\cancel{3y}-6v\cancel{-3y}\] Divide both sides by -6: \[\frac{u-3y}{-6}=\frac{\cancel{-6}v}{\cancel{-6}}\]\[v=\frac{u-3y}{-6}\]